# Understanding the Skill Gap in Recurrent Language Models: The Role of the Gather-and-Aggregate Mechanism

### v1

@@ 5 The G&A Bottleneck in SSMs
While both empirical evidence and theoretical analysis have established that individual SSMs struggle with retrieval tasks (Jelassi et al., 2024; Wen et al., 2024), this limitation has largely been examined in full-scale language models as a black-box phenomenon (Jelassi et al., 2024; Arora et al., 2023; Waleffe et al., 2024). Here, we provide a mechanistic account of how SSMs’ retrieval difficulties manifest within language models, focusing on the Gather-and-Aggregate mechanism implemented by just a few critical heads. Since the Transformer–SSM gap is most evident in retrieval-heavy tasks, this analysis supports the view that the gap arises not from broad architectural deficiencies, but from a small number of heads that fail to implement key algorithmic operations.
This section presents a sequence of findings that build toward this conclusion:
  • Section 5.1 visually inspects Aggregate Heads in Transformer and SSM models, showing that SSMs struggle with sharp, selective aggregation—likely due to their hidden-state architecture.
  • Section 5.2 shows that SSMs require more Aggregate Heads to match Transformer performance, suggesting the network compensates for the limited expressivity of individual heads through duplication.
  • Section 5.3 demonstrates that hybrid models delegate aggregation to attention layers. Disabling just six attention heads causes a 4×4$\times$4 × drop in MMLU, indicating that SSMs alone cannot support strong retrieval.
  • Section 5.4 identifies the Aggregate Heads as the critical bottleneck: replacing a single layer that contains them boosts MMLU from 33% to 50%, confirming their central role in retrieval.
  • Section 5.5 outlines strategies to overcome this bottleneck in SSM-based models, suggesting that targeted attention—either added during training or preserved during distillation—can restore G&A functionality.
@@ 5.1 Visual Inspection of G&A Heads
As shown in Figures 5 and 6 and discussed in Section 4.2, the Aggregate Heads in Mamba-1 and Mamba-2 correctly attend to the final tokens of each segment. However, they also smoothly attend to neighboring tokens, even though Section 4.3 shows that they can function without them. This likely introduces noise into the linear combination of segment summaries, reducing aggregation effectiveness. This smooth behavior aligns with hidden-state architectures, where information is compressed into a single representation over time, making it harder to capture the global nature of Aggregate Heads.
@@ 5.2 Redundancy of Aggregate Heads in SSM Models
    Gradual Degradation Reveals Redundant Retrieval Paths. As shown in Table 2 and discussed in Section 6, disabling heads in Llamba models leads to a slow, continuous drop in performance. This contrasts with the sharp collapses seen in Transformer-based models. The smoother degradation suggests that many SSM heads serve overlapping roles, indicating redundancy and less specialization.
@@
    Lack of Strong, Isolated Heads for Aggregation. Llamba-8B, distilled from Llama-3.1-8B using MOHAWK (Bick et al., 2024), mirrors the teacher’s structure by introducing G&A heads at layers 16 and 17. However, Table 7 shows a key difference: Llama-3.1-8B relies on two clearly effective heads (L17H24 and L17H27), while Llamba-8B depends on a single moderate head (L17H31), with others contributing minimally. Rather than isolating and concentrating aggregation in a few strong heads, Llamba spreads it across weaker alternatives.
@@
    More Heads Required to Match Performance. Llama-3.1-Minimal achieves 98% of its MMLU score with one Aggregate Head (Table 1). Llamba-8B-Minimal requires two to reach 95%, and Falcon-Mamba-7B-Minimal, built with lower-capacity Mamba-1 heads, needs four to reach 86% (Table 6). These differences indicate that SSM heads offer lower capacity individually and must be multiplied to maintain competitive performance.
@@ 5.3 Hybrid Models Delegate Aggregation to Attention
Attention layers in hybrid models are known to be critical for retrieval and copying tasks (Lieber et al., 2024; Glorioso et al., 2024; Waleffe et al., 2024). We hypothesize that during training, Aggregate Heads naturally concentrate in the attention layers, as attention’s greater expressivity makes it better suited for aggregation—underscoring a key limitation of SSMs.
To test this, we evaluate Zamba2-7B (Glorioso et al., 2024), the strongest 7B-scale hybrid model to date. Table 7 identifies attention heads critical for general retrieval; here, we further select nine that are especially important for MMLU: L47H{2, 4, 6, 8, 18, 21, 25} and L59H{17, 21}. Each head is disabled by zeroing out the final row of its attention matrix—a targeted intervention that impairs aggregation while leaving other capabilities, such as gathering and contextual mixing, largely intact. All SSM layers remain unchanged, preserving any aggregation they may contribute.
This intervention reduces MMLU accuracy from 64.3% to 34.9%, while knowledge task accuracy remains stable at 70%. Since MMLU’s random baseline is  25%, this 4x drop demonstrates that SSMs alone cannot sustain strong aggregation—hybrid models depend on attention layers to implement the G&A mechanism.
@@ 5.4 Hybrid Replacements
To further investigate whether SSMs struggle to implement the Aggregate Head, we examine how well an SSM mixer can replicate the functionality of an attention layer. We use the Llamba-8B (Bick et al., 2025) architecture, distilling each layer independently from Llama-3.1-8B (following Bick et al. (2024)) without end-to-end alignment. Specifically, the $i$-th Mamba-2 block in Llamba receives the output of Llama-3.1-8B’s $(i-1)$-th layer and is aligned independently using L2 distance against the output of Llama-3.1-8B’s $i$-th layer.
@@
The results, shown in Figure 7, reveal that most layer replacements had little or even a negative impact on MMLU. However, swapping L17, where we previously identified prominent Aggregate Heads at H24 and H27 (Table 1), significantly improved MMLU performance from 33% to 50%. To isolate their effect, we ran additional hybrid experiments: swapping only L17H24 while keeping the rest of the SSM heads intact increased performance to 44%, and including L17H27 raised it to 50%. This confirms that the contribution stems from the Aggregate Heads within L17, and further supports that individual SSM layers struggle to implement the Aggregate Head within G&A—reinforcing our claim that this limitation contributes to the performance gap between Transformers and SSMs.
@@ 5.5 Addressing G&A Limitations in SSM-Based Models
Efforts to improve the Gather-and-Aggregate mechanism within pure SSMs face a fundamental limitation: retrieval struggles stem from the fixed-size memory of RNN-style architectures, as shown in prior work (Jelassi et al., 2024; Wen et al., 2024). Rather than trying to overcome this constraint within the SSM design space, we explore hybrid models that incorporate attention layers to directly address these bottlenecks.
Our analysis reveals two empirical patterns that guide hybrid model design—one for training from scratch, and one for distillation:
@@
    Training hybrids from scratch. G&A heads consistently emerge in the middle of the network—for example, at layers L16 and L17 out of 32 in Llama-3.1-8B and Llamba-8B, and L35 and L36 out of 64 in Falcon-Mamba-8B. This suggests that attention is most valuable in mid layers, where retrieval operations naturally occur. As a result, when designing hybrid models from scratch, placing attention layers near the middle offers a targeted way to support G&A without relying on full attention throughout the network.
@@
These observations provide a practical recipe for hybrid design: introduce attention selectively—either guided by where G&A heads naturally emerge when training from scratch, or retained only in key locations when distilling from a Transformer. We leave further investigation into optimal placement and architectural variants of hybrid models to future work.
@@ 6 Transformer-SSM Gap Beyond MMLU
We previously showed that both Transformer- and SSM-based language models rely on a small number of heads to implement a two-stage Gather-and-Aggregate mechanism that supports in-context retrieval. This section extends that analysis to a broader range of benchmarks and model families, showing that differences in retrieval performance—and the gap between SSMs and Transformers in particular—often stem from specific algorithmic limitations, rather than from general deficiencies in language modeling ability.
We approach this in three stages: first, we isolate and characterize G&A heads across architectures; second, we quantify how architectural differences in G&A implementation affect task performance; and finally, we show how these differences are amplified or diminished by variations in task format.
@@ 6.1 Isolating Retrieval Differences Across Architectures
In Section 3, we showed that MMLU performance can be severely degraded by removing a single critical head—even when the model performs poorly on knowledge-focused tasks. This scenario is specific to MMLU: a task where strong retrieval alone can yield high accuracy, even when the model’s knowledge is degraded. In contrast, most benchmarks require both retrieval and factual knowledge, making it harder to isolate the contribution of retrieval alone.
@@
As described in Section 3.5, the synthetic KV-Retrieval task allows us to identify heads involved in G&A by measuring the impact of ablating each head on retrieval accuracy (Table 7). A significant drop in performance indicates that the head plays a role in the G&A mechanism. A similar approach was also used by Wu et al. (2024) to identify retrieval heads in Transformers. Crucially, because KV-Retrieval is synthetic and does not require external knowledge, it isolates retrieval ability without interference from knowledge degradation. This lets us trace G&A heads across layers and model families in a controlled and generalizable way.
@@ 6.2 Quantifying the Retrieval Gap
Having identified the retrieval-critical heads, we now measure their practical impact across benchmarks. Specifically, we incrementally disable the identified G&A heads in three model families:
  • Transformers: Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Touvron et al., 2023), which serve as baselines for effective G&A implementation.
  • Pure SSMs: Llamba-3B and Llamba-8B (Bick et al., 2025), built exclusively with state-space layers based on Mamba-2.
  • Hybrids: Zamba2-2.7B and Zamba2-7B (Glorioso et al., 2024), architectures primarily based on SSM layers supplemented by shared attention modules.
Table 2 summarizes the results.
In Zamba2 hybrid models, we ablate only attention heads while leaving SSM layers intact. As shown in Section 5.3, these models primarily perform retrieval via the attention layers, making them the relevant target for G&A ablation. Additionally, because attention is shared across layers in Zamba2, we disable just one instance of each shared head at a time, treating it as non-shared to isolate its effect.
In Transformer models, we observe a steep decline in performance on retrieval-heavy tasks after disabling only 10–20 heads, while accuracy on knowledge-focused tasks remains stable. This confirms that retrieval depends on a small, specialized set of heads. Zamba2-7B exhibits a similar pattern, albeit with nuances due to its shared-attention structure. Notably, different occurrences of the same shared head behave similarly when ablated, consistent with shared G&A roles, as seen in Table 7.
In contrast, Llamba-8B degrades gradually. Even after removing 30 G&A heads, it retains over 95% of its original performance across both retrieval and knowledge tasks. This slow decline reflects redundancy: SSM-based models rely on many weaker heads to approximate the function of fewer, sharper Transformer heads—aligning with our observations from Section 5.
@@ 6.3 Amplifying the Retrieval Gap through Task Format
Beyond architectural differences, the retrieval gap between Transformers and SSMs can be further amplified or mitigated by the task format itself. We previously saw hints of this in MMLU’s evaluation format, where rather than scoring the likelihood of full-text answer candidates, models must select a letter label (A, B, C, or D) corresponding to the correct option. To illustrate this effect, we compare two task variants with identical underlying content but different formats: ARC-Challenge in its original multiple-choice form versus ARC-Challenge Chat, a conversational variant.
As Table 3 shows, disabling G&A heads has minimal effect on the multiple-choice ARC-Challenge but drastically reduces accuracy in ARC-Challenge Chat. The conversational format thus increases retrieval demands by distributing relevant context throughout the dialogue, directly benefiting architectures with sharp retrieval mechanisms (Transformers and hybrids) over architectures with diffuse retrieval implementations (pure SSMs).
In this subsection, we examine a more subtle source of variation: the format of the task. Specifically, we compare two versions of ARC-Challenge that contain the same underlying questions but differ in presentation: the original multiple-choice format (used in our knowledge benchmarks), and a chat-style variant (ARC-Challenge Chat) shown in Table 3. While both versions rely on identical factual information, they respond very differently to G&A ablation.
ARC-Challenge Chat exhibits sharp performance drops—similar to MMLU—suggesting that its conversational structure introduces retrieval demands not present in the standard format. This highlights a key insight: retrieval load is shaped not only by what a task tests, but also by how it presents the information. In-context retrieval becomes more essential when relevant content is distributed across a dialogue, even if the underlying knowledge remains unchanged.
@@ 7 Conclusions
In this work, we show how both Transformer and SSM models develop Gather-and-Aggregate (G&A) mechanisms to support in-context retrieval, which in turn shapes performance on many popular benchmarks. While SSMs struggle to implement effective Aggregate Heads due to their fixed-size hidden state, this does not render them ineffective for language modeling.
We find that end-to-end models rely on only a small number of G&A heads for retrieval, while most other heads—supporting abilities such as parametric knowledge extraction and language understanding—perform comparably across architectures. This suggests that the overall Transformer-SSM performance gap may stem from specific algorithmic limitations confined to a few heads.
Hybrid models provide further insight: they naturally delegate the G&A mechanism to attention layers, leveraging their stronger algorithmic capacity. This explains why even a small number of attention layers can significantly enhance SSM-based models, and offers practical guidance for designing future hybrid architectures.
Beyond improving our understanding of sequence model mechanisms, our findings also offer insights into popular benchmarks and the distinction between knowledge and algorithmic skills. While MMLU is often regarded as a gold-standard benchmark that requires broad knowledge, our results indicate that MMLU performance is primarily governed by retrieval ability. This is evident from our finding that a 4B model, despite underperforming on knowledge-focused benchmarks, can still achieve strong MMLU performance (66%). This suggests that algorithmic skills—particularly retrieval—rather than pure knowledge retention, play a dominant role in MMLU success.
While this work highlights the role of G&A in both attention and SSM models, important open questions remain. Narrowing the retrieval gap is a step forward, but in-context learning and copying mechanisms warrant further investigation, as they may also relate to G&A. Furthermore, our hybrid model analysis focused on a limited set of architectures; a broader study could refine the optimal balance of attention and SSM layers.
@@ A.3 Performance in Minimal Models
A.4 Critical Heads in MMLU Performance
As highlighted in Section 3.4, we identify specific heads within the critical layers that are responsible for maintaining high MMLU scores. Table 1 presents data on the performance of Llama-3.1-8B-Minimal when different configurations of heads are retained. These results underscore the significance of certain heads, where even a single mixing head can carry enough information to achieve a near state-of-the-art MMLU score. The removal of these heads leads to a drastic drop in MMLU performance, confirming their essential role in retrieval tasks.
These results are complemented by the data for Falcon-Mamba-8B and Llamba-8B in Table 6 and Table 5, which reveal similar trends of critical head importance across different model architectures.
Heads Kept in a Layer Metrics (%)
0-15 16 17 MMLU Knowledge Tasks
0,1,…,31 22 24 66.32 39.09
0,1,…,31 $\emptyset$ 24 24.36 39.18
0,1,…,31 22 $\emptyset$ 25.59 39.21
0,1,…,31 $\emptyset$ $\emptyset$ 25.56 39.21
Table 4: Evaluating the Role of Critical Heads in Llama-3.1-8B-Minimal’s Performance. This table presents the impact of two critical attention heads (L16H22 and L17H25) on MMLU performance while keeping all preceding layers intact. Four configurations were tested: (1) both heads retained, (2) only the last-layer head retained, (3) only the second-to-last-layer head retained, and (4) neither head retained. Results show that these heads must work together to sustain high MMLU performance ($>$66%), while removing either or both drops accuracy to random guessing (25%). Note: This is a duplicate of Table 1, reprinted here for convenience in comparison with Tables 6 and 5.
Heads Kept in a Layer Metrics (%)
0-15 16 17 MMLU Knowledge Tasks
0,1,…,31 10 3,9 58.92 39.08
0,1,…,31 $\emptyset$ 3,9 42.17 39.16
0,1,…,31 10 $\emptyset$ 38.02 39.0
0,1,…,31 no L17MLP $\emptyset$ 29.82 36.76
Table 5: Performance evaluation of Llamba-8B-Minimal, focusing on the impact of three critical heads identified in L16H10, L17H3, and L17H9. The experiment examines MMLU performance under four configurations: (1) no critical heads retained, (2) only the 22nd head retained, (3) only the 24th head retained, and (4) both heads retained. The findings demonstrate that the coordination of these two heads is crucial for maintaining high MMLU performance, achieving over 58%. Removing either head results in a significant performance decline of nearly 20%. Additionally, eliminating the MLP learned through these heads causes a further reduction of approximately 10%. This additional drop is likely attributable to the presence of another G&A mechanism in the shallower layers.
Heads Kept in a Layer Metrics (%)
0-34 35 36 MMLU Knowledge Tasks
0,…,8191 3186, 5143 6607, 7305 1565, 1906 3873, 4925 52.26 37.23
0,…,8191 $\emptyset$ 1565, 1906 3873, 4925 35.91 37.25
0,…,8191 3186, 5143 6607, 7305 $\emptyset$ 39.39 37.26
0,…,8191 $\emptyset$ $\emptyset$ 24.98 37.28
Table 6: Performance evaluation of Falcon-Mamba-8B-Minimal under various configurations involving critical mixing heads. The experiment isolates four specific heads identified in the last two layers and evaluates MMLU scores in the following configurations: (1) all critical heads removed, (2) critical heads from the first layer retained, (3) critical heads from the second layer retained, and (4) all critical heads retained. The results reveal that only four heads, out of thousands at every layer are essential for leveraging the network’s knowledge to achieve an MMLU score of 52.26%. Excluding these heads results in a sharp performance drop to 25%, illustrating their critical role in transforming distributed knowledge into task-specific capabilities.
A.5 Retrieval Heads Across Model Families

### v2

@@ 5 The G&A Bottleneck in SSMs
The previous section established that both Transformers and SSMs rely on a small number of heads to implement a two-stage Gather-and-Aggregate (G&A) mechanism. We now examine how this mechanism becomes less effective in SSMs by focusing on the Aggregate stage, providing a mechanistic account of their retrieval limitations.
We begin by visually inspecting G&A behavior in SSMs (Section 5.1), showing that Aggregate Heads attend broadly rather than selectively. We then quantify redundancy in SSM aggregation (Section 5.2), finding that multiple heads are needed to match the performance of fewer, stronger Transformer heads. Next, we show that hybrid models offload aggregation to attention layers (Section 5.3); ablating just six heads causes a 4×4$\times$4 × drop in MMLU. We then identify these heads as the bottleneck (Section 5.4): replacing a single SSM layer restores MMLU from 33% to 50%. Finally, we explore how targeted attention during training or distillation can restore G&A in SSMs (Section 5.5).
@@ 5.1 Visual Inspection of G&A Heads
Figures 5 and 6 show that Mamba Aggregate Heads correctly attend to segment-final tokens but also attend unnecessarily to neighboring tokens. This likely introduces noise and reduces aggregation effectiveness by diluting focus on the intended summary tokens. The resulting smooth attention map reflects the constraints of hidden-state architectures, where temporal compression makes it harder to express the global behavior characteristic of Aggregate Heads.
@@ 5.2 Redundancy of Aggregate Heads in SSM Models
    Gradual Performance Decay. As shown in Table 2 and discussed in Section 6, ablating heads in Llamba causes a slow, steady drop in performance—unlike the sharp drops seen in Transformers—implying overlapping functionality and less specialization.
@@
    Absence of Strong Isolated Heads. Llamba-8B, distilled from Llama-3.1-8B using MOHAWK (Bick et al., 2024), introduces G&A heads at the same layers as the teacher (L16 and L17). However, Table 5 shows a key difference: Llama-3.1-8B relies on two clearly effective heads (L17H24 and L17H27), while Llamba-8B lacks strong, isolated heads. Its strongest (L17H31) is only moderately effective, and the others show minimal impact. Instead of concentrating aggregation in a few heads, Llamba distributes it across weaker ones—consistent with limited per-head expressivity.
@@
    Increased Head Count. Llama-3.1-Minimal achieves 98% of its MMLU score with one Aggregate Head. Llamba-8B-Minimal requires two to reach 95%, and Falcon-Mamba-7B-Minimal, built with lower-capacity Mamba-1 heads, needs four to reach 86% (see Table 4). These differences indicate that SSM heads offer lower capacity individually and must be multiplied to maintain competitive performance.
@@ 5.3 Hybrid Models Delegate Aggregation to Attention
Attention layers in hybrid models are known to be critical for retrieval and copying tasks (Lieber et al., 2024; Glorioso et al., 2024; Waleffe et al., 2024). We hypothesize that Aggregate Heads naturally concentrate in the attention layers during training, as attention’s greater expressivity makes it better suited for aggregation, underscoring a key limitation of SSMs.
To test this, we evaluate Zamba2-7B (Glorioso et al., 2024), the strongest 7B-scale hybrid model to date. Table 5 identifies attention heads critical for general retrieval; here, we further select nine that are especially important for MMLU: L47H{2, 4, 6, 8, 18, 21, 25} and L59H{17, 21}. Each head is disabled by zeroing out the final row of its attention matrix—impairing aggregation while preserving other capabilities like gathering and contextual mixing. All SSM layers remain unchanged, preserving any aggregation they may contribute.
This intervention reduces MMLU accuracy from 64.3% to 34.9%, while knowledge task accuracy remains stable at 70%. This 4×4$\times$4 × drop (relative to the 25% random baseline) demonstrates that hybrid models depend on attention layers for effective G&A implementation.
@@ 5.4 Hybrid Replacements
To further investigate whether SSMs struggle to implement the Aggregate Head, we examine how well an SSM mixer can replicate the functionality of an attention layer. We use the Llamba-8B (Bick et al., 2025) architecture, distilling each layer independently from Llama-3.1-8B (following Bick et al. (2024)) without end-to-end alignment. Specifically, the $i$-th Mamba-2 block in Llamba receives the output of Llama-3.1-8B’s $(i-1)$-th layer and is aligned using L2 distance against the $i$-th layer’s output.
@@
The results, shown in Figure 7, reveal that most layer replacements had little or even a negative impact on MMLU. However, replacing L17, which contains two prominent G&A heads at H24 and H27 (L17H24 was identified in Table 1), significantly improved performance from 33% to 50%. To isolate their effect, we ran additional hybrid experiments. Swapping only L17H24, while keeping the rest of the SSM heads intact, raised performance to 44%. Including L17H27, another G&A head not used in Table 1, increased performance to 50%. This confirms that the contribution stems from the Aggregate Heads within L17, and further supports that individual SSM layers struggle to implement the Aggregate Head within G&A—reinforcing our claim that this limitation contributes to the performance gap between Transformers and SSMs.
@@ 5.5 Addressing G&A Limitations in SSM-Based Models
Efforts to improve the Gather-and-Aggregate mechanism within pure SSMs face a fundamental limitation: retrieval struggles stem from the fixed-size memory of RNN-style architectures, as shown in prior work (Jelassi et al., 2024; Wen et al., 2024). Rather than trying to overcome this constraint within the SSM design space, we explore hybrid models that incorporate attention layers to directly address these bottlenecks. Our analysis reveals two empirical patterns that guide hybrid model design—one for training from scratch, and one for distillation:
@@
    Training hybrids from scratch. G&A heads consistently emerge in the middle of the network—for example, at layers L16 and L17 out of 32 in Llama-3.1-8B and Llamba-8B, and L35 and L36 out of 64 in Falcon-Mamba-7B. This suggests that attention is most valuable in mid layers, where retrieval operations naturally occur. As a result, when designing hybrid models from scratch, placing attention layers near the middle offers a targeted way to support G&A without relying on full attention throughout the network.
@@
These findings suggest a simple recipe for hybrid design: insert attention where G&A heads emerge during training, or preserve it at key layers when distilling from a Transformer. Determining optimal placement and refining architectural choices remain open directions for future work.
@@ 6 Transformer-SSM Gap Beyond MMLU
We previously showed that both Transformers and SSMs rely on a few heads to implement a two-stage Gather-and-Aggregate (G&A) mechanism for in-context retrieval. We now extend this analysis to additional benchmarks and models, reinforcing that the broader retrieval gap stems from specific algorithmic limits rather than general language modeling ability. Our analysis has three parts: Section 6.1 identifies G&A heads across models; Section 6.2 evaluates how architectural choices affect performance; Section 6.3 examines how task format influences these effects.
@@ 6.1 Isolating Retrieval Differences Across Architectures
In Section 3, we showed that MMLU performance can be severely degraded by removing a single critical head—even when the model performs poorly on knowledge-focused tasks. This scenario is specific to MMLU, where strong retrieval alone can yield high accuracy even with degraded knowledge. In contrast, most benchmarks require both retrieval and factual knowledge, making it harder to isolate the contribution of retrieval alone.
@@
As described in Section 3.5, the synthetic KV-Retrieval task allows us to identify heads involved in G&A by measuring the impact of ablating each head on retrieval accuracy (Table 5). A significant drop in performance indicates the head contributes to the G&A mechanism. A similar approach was also used by Wu et al. (2024) to identify retrieval heads in Transformers. Crucially, since KV-Retrieval is synthetic and requires no external knowledge, it isolates retrieval ability without confounding effects from knowledge loss. This lets us trace G&A heads across layers and model families in a controlled and generalizable way.
@@ 6.2 Quantifying the Retrieval Gap
Having identified the retrieval-critical heads, we now assess their practical impact across benchmarks by incrementally disabling them in three model families. These include Transformer models (Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Touvron et al., 2023)) as baselines for effective G&A implementation; pure SSM models (Llamba-3B and Llamba-8B (Bick et al., 2025)) based on Mamba-2; and hybrid models (Zamba2-2.7B-Instruct and Zamba2-7B-Instruct (Glorioso et al., 2024)) that combine Mamba-2 layers with shared attention modules. In Zamba2, we ablate only attention heads, as they are responsible for retrieval (Section 5.3); each shared head is ablated once to isolate its effect. For clarity, we omit instructional and embedding-tying suffixes from all model names.
Table 2 summarizes the results. In Transformer models, we observe a steep decline in performance on retrieval-heavy tasks after disabling only 10–20 heads, while accuracy on knowledge-focused tasks remains stable. This confirms that retrieval depends on a small, specialized set of heads. Zamba2-7B exhibits a similar pattern, albeit with nuances due to its shared-attention structure. Notably, different occurrences of the same shared head behave similarly when ablated, consistent with shared G&A roles, as seen in Table 5.
In contrast, Llamba-8B degrades gradually. Even after removing 30 G&A heads, it retains over 95% of its original performance across both retrieval and knowledge tasks. This slow decline reflects redundancy: SSM-based models use many weaker heads to approximate the function of fewer, sharper Transformer heads—aligning with our observations from Section 5.
@@ 6.3 Amplifying the Retrieval Gap through Task Format
Beyond architecture, task format can amplify or reduce the retrieval gap between Transformers and SSMs. We previously noted this in MMLU, where models select a letter label rather than score full answer texts.
To further illustrate, we compare two formats of the same content: the original ARC-Challenge multiple-choice task and its conversational variant, ARC-Challenge Chat. As shown in Table 3, disabling G&A heads has little effect on the multiple-choice version but sharply reduces accuracy in the chat variant. The conversational format increases retrieval demands by spreading relevant content across dialogue turns, favoring architectures with sharper retrieval (Transformers and hybrids) over those with more diffuse mechanisms (pure SSMs).
This highlights a key point: retrieval pressure depends not just on what is asked, but on how it’s presented. When context is more distributed, effective in-context retrieval becomes more critical—even if the underlying knowledge stays the same.
@@ 7 Conclusions
We show that both Transformer and SSM models develop a Gather-and-Aggregate (G&A) mechanism across a small number of heads to support in-context retrieval, which is crucial for performance on many benchmarks. While SSMs struggle to implement strong G&A heads due to their fixed-size hidden state, this limitation is localized and does not reflect a general weakness in language modeling.
Hybrid models provide further insight: they naturally offload G&A to attention layers, leveraging their stronger algorithmic capacity. This explains why even a few attention layers can meaningfully boost SSM models, providing practical guidance for hybrid design.
Our results also reframe benchmark interpretation. Although MMLU is often seen as a broad knowledge test, we find that MMLU performance is primarily governed by retrieval ability. This highlights the role of algorithmic skills, and particularly retrieval, in benchmark success.
Important open questions remain: while narrowing the retrieval gap is a key step, the role of G&A in in-context learning and copying warrants further study. It also remains to be seen whether G&A is part of a broader retrieval process involving other head types, such as Amplification Heads (Lieberum et al., 2023), which enhance signal propagation.
@@ A.3 Performance in Minimal Models
A.4 Critical Heads Drive MMLU Performance Across Models
To assess how in-context retrieval is localized within specific heads, we examine the MMLU performance of three models—Llama-3.1-8B, Llamba-8B, and Falcon-Mamba-7B—under targeted head ablation. Table 4 presents the results of selectively retaining or removing mixing heads in the final two layers of each model, while preserving all earlier components.
Across all architectures, we observe a sharp and consistent pattern: removing just one or two critical heads reduces MMLU accuracy to near-random levels ($\sim$25%), while retaining them restores much of the original performance. Notably, these heads represent fewer than 0.1% of the total head count, highlighting the extreme concentration of retrieval ability within a tiny subset of the model. Meanwhile, scores on knowledge-focused benchmarks remain nearly unchanged across all configurations, reinforcing that these heads are specifically responsible for in-context retrieval rather than general language modeling or factual recall.
This result supports the broader claim that the Gather-and-Aggregate mechanism emerges in just a handful of components and suggests that hybrid and efficient model designs can selectively preserve only these heads.
Model Heads Retained Metrics (%)
Prev. Layers Layer X Layer X+1 MMLU Knowledge Tasks
Llama-3.1-8B 0–31 22 24 66.32 39.09
0–31 $\emptyset$ 24 24.36 39.18
0–31 22 $\emptyset$ 25.59 39.21
0–31 $\emptyset$ $\emptyset$ 25.56 39.21
Llamba-8B 0–31 10 3, 9 58.92 39.08
0–31 $\emptyset$ 3, 9 42.17 39.16
0–31 10 $\emptyset$ 38.02 39.00
0–31 $\emptyset$ $\emptyset$ 29.82 36.76
Falcon-Mamba-7B 0–8191 3186, 5143 6607, 7305 1565, 1906 3873, 4925 52.26 37.23
0–8191 $\emptyset$ 1565, 1906 3873, 4925 35.91 37.25
0–8191 3186, 5143 6607, 7305 $\emptyset$ 39.39 37.26
0–8191 $\emptyset$ $\emptyset$ 24.98 37.28
Table 4: Impact of Critical Head Removal Across Model Architectures. We evaluate how retaining or removing specific heads in the final two layers affects MMLU and knowledge task performance for three models: Llama-3.1-8B, Llamba-8B, and Falcon-Mamba-7B. Each row shows performance under a different configuration, where the ”Heads Retained” columns indicate which heads are active in the two final layers (”Layer X” and ”Layer X+1”) while all earlier heads are preserved. The heads shown correspond to the critical Gather and Aggregate Heads identified in prior analyses. For Llama and Llamba models, retaining both heads restores near-original MMLU performance (66.3% and 58.9%, respectively), while removing either or both causes performance to collapse to near-random levels. In Falcon-Mamba-7B, just four attention channels (among thousands) across the final two layers are sufficient to recover 52.3% MMLU accuracy—again, comparable to full-model performance. Knowledge task scores remain mostly unchanged across configurations, reinforcing that these heads specifically control retrieval rather than factual knowledge access. This shows that across architectures, in-context retrieval is driven by just a few heads, whose removal alone accounts for the majority of the performance gap. Note: The information for Llama-3.1-8B is a duplicate of Table 1, reprinted here for convenience in comparison.
A.5 KV-Retrieval Recipe
To probe retrieval behavior in isolation, we design a synthetic memorization task in which the model is asked to recall values associated with keys presented earlier in the same prompt. Each input contains a list of key-value (KV) pairs, followed by a query for the value of one of the keys. For example:
Memorize the following dictionary:
present:50
institute:0
scallops:84
neuropsychiatry:67
The value of the key ’scallops’ is
Evaluation Setup.
We use dictionaries containing key-value pairs and evaluate the model on randomly generated examples. For each example, we assess the model’s ability to retrieve the correct value under two settings:
  1. 1.
    Generation: The model is prompted with the full dictionary and asked to generate 10 additional tokens. We check whether the correct value appears anywhere in the generated output.
  2. 2.
    Answer Scoring: Each possible value from the dictionary is appended to the end of the prompt, and the log-probability of the resulting sequence is computed. The value assigned the highest log-probability is treated as the model’s prediction.
The key difference between these strategies lies in how the model expresses its beliefs. Generation allows the model to unfold a short sequence of reasoning steps, operating within a richer space of token-level interactions. However, this flexibility comes at the cost of interpretability, making it harder to surgically probe the model’s internal mechanisms. Answer scoring, by contrast, restricts the model to a fixed set of candidates, offering a more direct readout of its internal ranking.
This distinction is important because generation can surface hallucinations—cases where the model outputs values not present in the dictionary or otherwise digresses from the task—while answer scoring is restricted to valid options. We find that the heads responsible for hallucinations differ from those that enable correct retrieval. This stands in contrast to the formulation of Wu et al. [2024], who used generation alone and concluded that “retrieval heads” may be responsible for hallucinations.
Prompt Format.
We test two variants of the query prompt that differ only in their final character:
  1. 1.
    ‘The value of the key ‘scallops’ is‘ and
  2. 2.
    ‘The value of the key ‘scallops’ is ‘ (with a trailing space).
Though superficially similar, these variants elicit markedly different behaviors under Answer Scoring. In variant (1), the model must evaluate completions like ‘The value of the key ‘scallops’ is84‘, which lacks a space and results in unnatural tokenization. Surprisingly, this unnatural format is highly diagnostic: head-level retrieval signals emerge clearly with as few as 20–30 key-value pairs. In contrast, variant (2), which uses the more natural space-delimited form, requires much longer dictionaries to produce similar head-level effects. For comparison, Wu et al. [2024] relied on 35K-token contexts to observe such patterns—well beyond typical context lengths.
We hypothesize that this discrepancy reflects a shift in the distribution of retrieval effort across the model. When the prompt ends cleanly on a token boundary, the model can offload the task to a small number of dedicated retrieval heads. By contrast, introducing a continuation token (like a space) appears to encourage more distributed processing across heads.
Head-Ablation Protocol.
To identify attention heads involved in retrieval, we systematically ablate each head by zeroing out its output projection and measure the resulting drop in accuracy under Answer Scoring (without the trailing space). This procedure isolates heads that are critical for retrieval while keeping computational overhead low.
Overall, this setup offers a clean and flexible diagnostic for in-context retrieval—independent of natural language or factual knowledge—and clarifies how individual heads contribute to this core algorithmic skill.
A.6 Retrieval Heads Across Model Families